# Preprint Manuscript

### v2

@@
Second Version
CUBIC INCOMPLETENESS: HILBERT’S TENTH PROBLEM BEGINS AT DEGREE THREE
@@
(September 2025)
We resolve a long-standing open question in the affirmative: cubic Diophantine equations are undecidable. By encoding Gödel’s incompleteness theorem via Fibonacci-based Gödel numbering and Zeckendorf representation, combined with a systematic guard-gadget framework, we construct an explicit degree-3 polynomial $P_{\text{hard}}\in\mathbb{Z}[x_{1},\ldots,x_{M}]$ whose solvability over $\mathbb{N}$ is equivalent to provability of a Gödel sentence in Peano Arithmetic. We establish three results. First, a bidirectional reduction proving $P_{\text{hard}}(\vec{x})=0$ has solutions if and only if $G_{T}$ is provable. Second, a uniform construction showing that for any Turing machine $M$, we can effectively produce a cubic polynomial $Q_{M}(\vec{u})\in\mathbb{Z}[\vec{u}]$ such that $Q_{M}(\vec{u})=0$ has a solution in $\mathbb{N}^{k}$ if and only if $M$ halts, establishing undecidability of the class $\delta=3$. Third, we formalize an impredicativity thesis: no consistent theory extending Robinson arithmetic can define the minimal degree at which Diophantine undecidability emerges.
@@ 1 Overview
1.1 Corrigendum
The present revision stands as a clarification of intent and method. What remains unchanged is the theoretical vision; what has been refined is the machinery by which that vision finds expression. We now provide a systematic theoretical enumeration of all polynomial degrees throughout the encoding. Critically, the earlier version contained a fatal error: the result of the “promissory” shielding process was incorrectly claimed to be of degree three. In this revision, we omit the problematic equation and focus instead on explicit witness construction, ensuring all constraints remain genuinely cubic.
1.2 Introduction
Previously [^1] A Fibonacci-Based Gödel Numbering: $\Delta_{0}$ Semantics Without Exponentiation, [2025], arXiv:2509.10382 [math.LO], https://arxiv.org/abs/2509.10382, we proposed that a Zeckendorf-based arithmetization of Gödel numbering—eschewing exponentiation in favor of additive structure—might lower the formal overhead in encoding undecidability. We now apply this method to produce undecidable Diophantine equations of degree three, thereby advancing “Hilbert’s Tenth Problem” [^2] As posed by Hilbert [1900]. into the cubic regime ($\delta=3$) and resolving a previously open case: We cite from the article “Universal Diophantine Equation”, by Jones [1982]:
@@ 1.2 Introduction
Problem 1.1 (Cubic Remainder of Hilbert’s 10th Problem).
@@
$P(x_{1},\ldots,x_{n})=0,\quad P\in\mathbb{Z}[x_{1},\ldots,x_{n}],\quad\delta(P)\leq 3.$ (1.1)
@@
1.3 Contribution
We construct a witness of polynomial Diophantine equation of total degree at most three, whose solvability over $\mathbb{N}$ is equivalent to the provability of a Gödel sentence in Peano Arithmetic extended with Fibonacci-based encoding. This resolves the question of whether undecidability can be exhibited at cubic degree, thereby advancing the study of Hilbert’s Tenth Problem into the low-degree regime.
Theorem 1.2 (Main Result).
There exists a polynomial $P_{\text{hard}}\in\mathbb{Z}[x_{1},\ldots,x_{M}]$ of $\delta(P_{\text{hard}})\leq 3$ and a sentence $G_{T}$ in the language of arithmetic such that:
$\exists\vec{x}\in\mathbb{N}^{M}\colon P_{\text{hard}}(\vec{x})=0\quad\Longleftrightarrow\quad T_{\text{Fib}}\vdash G_{T},$ (1.3)
where $T_{\text{Fib}}$ denotes Peano Arithmetic with Fibonacci-based Gödel numbering, and $G_{T}$ is the canonical Gödel sentence asserting its own unprovability.
The logic encodes proof-theoretic relations—axiom membership, modus ponens, and sequence verification—into polynomial constraints of controlled degree. The key innovation is a guard-gadget framework that transforms potentially negative constraints into sums of squares, preserving Diophantine equivalence while maintaining cubic degree bounds. The results of Matiyasevich [1970] and Robinson et al. [1961] establishes that every recursively enumerable set can be represented by a Diophantine equation. Subsequent refinements by Jones [1982] proved realizability at $\delta\leq 4$. We build on this tradition in two key ways:
@@ 1.3 Contribution
    By Explicit proof verification: We forward the arithmic content of a formal proof (axioms, modus ponens) directly in $\Delta_{0}$, rather than constructing a $\Sigma_{1}$ abstraction to minimize overhead.
@@
    By Finite axiom sets: By working with a precalculated representations, we avoid the need to encode an infinite, or expressive axiom scheme.
The trade-off is that our polynomial is potentially “wider” (in number of variables and monomials) than a optimal $\Sigma_{1}$ construction, but the argument is arguably more elementary (e.g., in terms of complexity), relying primarily on a variant of Gödel’s $\beta$-function (Theorem 2.1) and Zeckendorf representation (Definition 2.3). Having outlined our approach, we now formalize the key technical preliminaries underpinning the construction.
@@ 2.1 The Gödel Beta-Function
Procedure 2.1 (Diophantine $\beta$-Function).
@@
In the completeness proof of Theorem 3.16, the witness construction algorithm (Exposition 3.9) must compute four-square decompositions for each remainder $r_{i}$ and its bound $(i+1)d-1-r_{i}$. By the above algorithm, this can be done in time $O(\log((i+1)d))$ per decomposition, yielding total time $O(N\cdot\log(\max_{i}f_{i}))$ for all $N$ proof lines. This is polynomial in the proof parameters, as required.
@@ 2.3 Combined Encoding
    Encode the sequence using $\beta(c,d,i)=f_{i}$ via constraint (3.34) and Lemma $2.$
@@ 2.4 Summary of Polynomial Constraints
Constraint Equation $n\leq\delta$
$\beta$-Division $c-q_{i}(1+(i+1)d)-r_{i}=0$ 2
Lower bound $r_{i}-a_{i}^{2}-b_{i}^{2}-c_{i}^{2}-e_{i}^{2}=0$ 2
Upper bound $(i+1)d-1-r_{i}-\alpha_{i}^{2}-\beta_{i}^{2}-\gamma_{i}^{2}-\delta_{i}^{2}=0$ 2
Boolean digits $d_{i,\kappa}-d_{i,\kappa}^{2}=0$ 2
Non-adjacency $d_{i,\kappa}\cdot d_{i,\kappa+1}=0$ 2
Zeckendorf sum $f_{i}-\sum_{\kappa=2}^{K}F_{\kappa}d_{i,\kappa}=0$ 1
Table 1: All constraints have degree ($\delta$) $\leq 2$ before applying guard multipliers (see Section 2.5).

### v3

@@
Third Version
CUBIC INCOMPLETENESS: HILBERT’S TENTH PROBLEM OVER $\mathbb{N}$ STARTS AT $\delta=3$
@@
(Oktober 2025)
We prove that Hilbert’s Tenth Problem over $\mathbb{N}$ remains undecidable when restricted to cubic equations (degree $\leq 3$), resolving the open case $\delta=3$ identified by Jones (1982) and establishing sharpness against the decidability barrier at $\delta=2$ (Lagrange’s four-square theorem). For any consistent, recursively axiomatizable theory $T$ with Gödel sentence $G_{T}$, we effectively construct a single polynomial $P(x_{1},\ldots,x_{m})\in\mathbb{Z}[\mathbf{x}]$ of degree $\leq 3$ such that $T\vdash G_{T}$ if and only if $\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0$. Our reduction proceeds through four stages with explicit degree and variable accounting. Proof-sequence encoding via Diophantine $\beta$-function and Zeckendorf representation yields $O(KN)$ quadratic constraints ($K=O(\log(\max_{i}f_{i}))$, NN = proof length). Guard-gadgets wrap each constraint $E(\mathbf{x})=0$ as $u\cdot E(\mathbf{x})=0$, $u=1+v^{2}$, preserving degree $\leq 3$ with $O(KN^{3})$ variables. Sum-of-squares aggregation $P_{\text{merged}}=\sum_{i}P_{i}^{2}$ produces degree $\leq 6$ with $O(KN^{3})$ monomials. Recursive monomial shielding reduces degree to $\leq 3$ in $O(\log d)$ rounds, adding $O(K^{3}N^{3})$ variables total.
@@
1 Corrigendum
1.1 First Corrigendum
The initial version encoded a Gödel sentence as a Diophantine system rather than a single equation. The ambition to present a concrete standalone equation introduced an unwarranted bias into the construction. Subsequent algorithmic approaches yielded a massive closed form, but the reduction diverted attention from the central epistemic questions.
1.2 Second Corrigendum
In the second revision, we pivoted to explicit witness construction, ensuring all constraints remain genuinely cubic, and grounded the argument in logical foundations rather than forcing a premature encoding.
1.3 Third Corrigendum
In this third—still preliminary—version, we resolve ambiguities between logical and mathematical layers ($\mathbb{Z}$ versus $\mathbb{N}$), and improve on bookkeeping. The first-order argument is developed independently of its Diophantine realization. We provide a high-level overview at the outset, then verify each layer on its own terms.
[…]
We now adopt a direct approach: all variables, constraints, and solutions are understood to inhabit $\mathbb{N}$ from the outset. The $\beta$-function (Theorem 3.1), four-squares encoding (Lemma 4.2), and Zeckendorf representation (Definition 3.3) operate naturally over $\mathbb{N}$. The guard-gadget framework (Definition 3.6) enforces positivity via $u=1+v^{2}\geq 1$ without recourse to negative integers. This interpretation aligns with the historical formulation of Hilbert’s Tenth Problem over $\mathbb{N}$ and avoids unnecessary detours through $\mathbb{Z}$. The approach is straightforward: Turing machines operate with discrete symbols.
1.4 Context
This paper shares ideas with different work [^1] The Fractal Logic of $\Phi$-adic Recursion, arXiv:2510.08934, [2025a], in that a compression-undecidability relationship is visible in the context of extensions from $\mathbb{N}$ to $\mathbb{Z}$. The additional compression offered by negative numbers permits lower polynomial degrees in Diophantine equations, but at the cost of increased uncertainty between powers. The equation $x^{2}=4$ is uniquely solved in $\mathbb{N}$ by $x=2$, while in $\mathbb{Z}$ the solution set comprises $\{-2,2\}$. More dramatically, $x^{4}=16$ admits four solutions $\{\pm 2,\pm 2i\}$ in $\mathbb{C}$, with the real solutions $\pm 2$ already exhibiting the doubling. Future work will focus on this: The ambiguity is not a deficiency but a necessary consequence of Kolmogorov [1968] complexity: syntactic reduction of degrees is purchased through semantic branching. Results possessing nontrivial structure in $\mathbb{N}$ collapse in $\mathbb{Z}$ to trivial symmetries, whereby the displaced complexity migrates into sign ambiguity.
@@ 2 Overview
2.1 The Problem
Previously [^2] A Fibonacci-Based Gödel Numbering: $\Delta_{0}$ Semantics Without Exponentiation, arXiv:2509.10382, [2025b], we proposed that a Zeckendorf-based arithmetization of Gödel numbering—eschewing exponentiation in favor of additive structure—might lower the formal overhead in encoding undecidability. We now apply this method to produce undecidable Diophantine equations of degree three, thereby advancing “Hilbert’s Tenth Problem” [^3] As posed by Hilbert [1900]. into the cubic regime ($\delta=3$) and resolving a previously open case: We cite from the article “Universal Diophantine Equation”, by Jones [1982]:
@@ 2.1 The Problem
Problem 2.1 (Cubic Remainder of Hilbert’s 10th Problem).
@@
$P(x_{1},\ldots,x_{n})=0,\quad P\in\mathbb{N}[x_{1},\ldots,x_{n}],\quad\delta(P)\leq 3.$ (2.1)
@@
2.2 Roadmap
[…]
The restriction to $\mathbb{N}$ rather than $\mathbb{Z}$ is essential. Cubic equations over $\mathbb{Z}$ admit decidability through Thue equations (Baker’s transcendence methods), Hasse principle obstructions, and elliptic curve descent via Mordell–Weil theory. Undecidability over $\mathbb{N}$ exploits combinatorial complexity introduced by non-negativity constraints, enforced polynomially through guard-gadgets: replacing constraints $E(\mathbf{x})=0$ of degree at most two with the guarded system $u\cdot E(\mathbf{x})=0$, $u=1+v^{2}$ maintains cubic degree while forcing $u\geq 1$. Naive activation variables or difference-of-squares encodings produce degree-four terms, necessitating this construction.
Stage Configuration Implication
Step 1 $\mathcal{A}_{1}$ exists $\delta=2$ decidable (four-squares)
Step 2 $\mathcal{A}_{1}\land\mathcal{A}_{2}$ exist Can reduce any $\delta=6$ to equivalent $\delta=3$
Step 3a Assume $\mathcal{A}_{3}$ exists Contradiction: $\mathcal{A}_{1}\land\mathcal{A}_{2}\land\mathcal{A}_{3}$ decides PA
Step 3b Reject hypothesis Conclusion: $\delta=3$ undecidable over $\mathbb{N}$
Table 1: Logical structure of the impossibility proof.
The technical realization is a witness of a single cubic polynomial $P_{\text{hard}}(\mathbf{x})$ with $m=O(K^{3}N^{3})$ variables, where $N$ is the proof length and $K=O(\log(\max_{i}f_{i}))$ bounds Zeckendorf digit count. Every intermediate constraint—after guard-wrapping—has degree $\leq 3$ (by explicit bookkeeping, see 2); sum-of-squares merging yields degree $\leq 6$, reduced to cubic via monomial shielding (Definition 6.1) with $O(K^{3}N^{3})$ auxiliary variables
$P_{\text{merged}}=\sum_{i}P_{i}^{2}\in\delta(6)$ (2.3)
followed by recursive monomial shielding: each monomial $x^{a}y^{b}z^{c}$ with $a+b+c>3$ factors as $p\cdot q$ with $\delta(p),\delta(q)\leq 3$, introducing auxiliary $w$ via $w=p\cdot q$. Iteration terminates in $O(1)$ rounds, yielding a single cubic polynomial $P_{\text{hard}}(\mathbf{x})$ with $O(K^{3}N^{3})$ variables. The variable inflation from systems to single equations reflects the cost of degree-preserving aggregation.
Our proof by contradiction states: No universal cubic equation $P_{\text{univ}}(n,\mathbf{x})$ can encode all halting problems: such a polynomial would render Chaitin [1975] constant
$\Omega=\sum_{M\text{ halts}}2^{-|M|}$ (2.4)
computable by enumeration. Our construction avoids this through non-uniformity: $P_{\text{hard}}$ depends on the uncomputable proof length $N$. This applies equally to systems and single equations, though systems admit sparse structure while single equations become dense under sum-of-squares.
[…]
The cubic barrier represents a genuine phase transition in Diophantine complexity over $\mathbb{N}$, arising from arithmetic structure, non-negativity, and computational irreducibility: systems achieve cubic degree through encoding design, while single equations require degree-reduction post-processing.
Figure 1: A high-level overview of the reduction from provability in Peano Arithmetic to the solvability of cubic Diophantine systems. The construction proceeds through seven stages, each maintaining polynomial bounds on variable count and equation degree.
2.2.1 Encoding Proof Sequences (Stages 1–3)
A proof $\pi=\langle\varphi_{1},\ldots,\varphi_{N}\rangle$ is encoded via Gödel numbers $f_{i}\in\mathbb{N}$. The sequence is stored using two complementary representations:
@@ 2.2.1 Encoding Proof Sequences (Stages 1–3)
    $\beta$-Function Storage (Theorem 3.1): Existentially quantify over parameters $c,d\in\mathbb{N}$ satisfying
    $\beta(c,d,i)=f_{i}\quad\text{for all }i\in\{1,\ldots,N\},$ (2.5)
    where $\beta(c,d,i)$ denotes the remainder when dividing $c$ by $1+(i+1)d$. This is enforced via quadratic constraints (Lemma 4.1):
    $c$ $=q_{i}\cdot(1+(i+1)d)+r_{i},$ (2.6)
[…]
@@
    Zeckendorf Digits (Definition 3.3): Each $f_{i}$ is represented as a sum of non-consecutive Fibonacci numbers:
    This representation is unique and enables arithmetic verification without carry propagation (Exposition Exposition).
2.2.2 Proof Verification (Stage 4)
Each line $i$ must be justified either as an axiom or by modus ponens from earlier lines. This disjunction is encoded via boolean activation variables:
  1. 1.
    Axiom Test (Lemma 4.5): For a finite axiom set $\{g_{1},\ldots,g_{M}\}$, the constraint
    $b_{ax,i}\cdot\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0,\quad b_{ax,i}\in\{0,1\}$ (2.9)
    forces $f_{i}\in\{g_{1},\ldots,g_{M}\}$ when $b_{ax,i}=1$ (since a sum of squares vanishes iff each term vanishes).
  2. 2.
    Modus Ponens Test (Lemma 4.7): For each triple $(i,j,k)$ with $j,k<i$, the constraint
    $b_{mp,i,j,k}\cdot(f_{i}-f_{j}-f_{k})=0,\quad b_{mp,i,j,k}\in\{0,1\}$ (2.10)
    forces $f_{i}=f_{j}+f_{k}$ when $b_{mp,i,j,k}=1$. The Zeckendorf digits of $f_{i}$ automatically satisfy the non-adjacency constraint by uniqueness (Remark Exposition).
  3. 3.
    Exactly-One Justification (Lemma 4.11): The constraint
    $b_{ax,i}+\sum_{j,k<i}b_{mp,i,j,k}=1$ (2.11)
    ensures that exactly one justification mechanism activates per line.
2.2.3 Degree Management (Stages 5–7)
The proof verification constraints (Stage 4) produce equations of degree $\leq 3$, but merging them into a single equation via sum-of-squares yields degree $6$. We employ two techniques:
  1. 1.
    Guard-Gadgets (Definition 3.6): Each constraint $E(\mathbf{x})=0$ with $\delta(E)\leq 2$ is replaced by the guarded system
    $u\cdot E(\mathbf{x})$ $=0,$ (2.12)
    $u-1-v^{2}$ $=0,$ (2.13)
    where $u,v\in\mathbb{N}$. The second equation forces $u=1+v^{2}\geq 1$, hence $u\neq 0$, so the first equation implies $E(\mathbf{x})=0$. This increases degree to at most $\max(1+2,2)=3$.
  2. 2.
    Jones Factorization (Theorem 5.1): After merging via sum-of-squares (yielding $P_{\text{merged}}$ of degree $6$), we apply recursive monomial shielding. Each monomial $x^{a}y^{b}z^{c}$ with $a+b+c>3$ is factored as $p\cdot q$ where $\delta(p),\delta(q)\leq 3$, introducing auxiliary variables $w$ via constraints $w=p\cdot q$ of degree $\leq 3$. This process (Algorithm 6.3) introduces $O(m\log d)$ variables for a polynomial with $m$ monomials of degree $\leq d$.
2.3 Main Result
Theorem 2.2 (Restatement of Theorem 5.2).
Let $\mathrm{T}$ be a recursively axiomatizable theory with finite axiom set $\mathcal{A}$, and let $G_{T}$ be its Gödel sentence. There exists a polynomial $P(\mathbf{x})\in\mathbb{Z}[\mathbf{x}]$ of degree $\leq 3$ such that:
The polynomial has $m=O(K^{3}N^{3})$ variables, where $N$ is the proof length and $K=O(\log(\max_{i}f_{i}))$ bounds the Fibonacci representation size.
2.4 On Single Cubic Equations
While our construction produces a system of cubic equations, one might ask whether a single cubic equation could suffice. The following obstruction shows this is impossible.
Proposition 2.3 (Impossibility of Universal Cubic Equation).
There exists no polynomial $P_{\text{univ}}(n,\mathbf{x})\in\mathbb{Z}[n,\mathbf{x}]$ within an $\omega$-consistent formal system of degree $\leq 3$ such that for all Turing machines $M$ (encoded as $n\in\mathbb{N}$):
$M\text{ halts}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{k}:P_{\text{univ}}(n,\mathbf{x})=0.$ (2.15)
Proof sketch.
Assume such $P_{\text{univ}}$ exists. For each machine $M$ encoded as $n$, enumerate solutions $\mathbf{x}\in\mathbb{N}^{k}$ in order of increasing norm $|\mathbf{x}|_{1}=\sum_{i}x_{i}$. By the degree-$3$ constraint, the polynomial $P_{\text{univ}}(n,\mathbf{x})$ has at most $O(k^{3})$ monomials, so checking whether $P_{\text{univ}}(n,\mathbf{x})=0$ takes polynomial time in $|\mathbf{x}|_{1}$.
For halting machines, the smallest solution $\mathbf{x}^{*}$ satisfies $|\mathbf{x}^{*}|_{1}=O(\text{poly}(n))$ (by the polynomial encoding of bounded computations). Thus, we could compute Chaitin [1975] constant.
$\Omega=\sum_{M\text{ halts}}2^{-|M|}$ (2.16)
by enumerating solutions up to polynomial size—contradicting the incomputability of $\Omega$, hence no such $P_{\text{univ}}$ exists. ∎
Remark.
Our construction avoids this obstruction by producing a machine-dependent system $\Phi_{M}(\mathbf{x})$. The proof length $N$ (and hence the variable count $m=O(K^{3}N^{3})$) is not computable from the machine description alone—this non-uniformity is essential. The system encodes "there exists a proof of length NN", not "the shortest proof has length NN".
2.5 Prior Work and Roadmap
We recall:
  1. 1.
    Robinson et al. [1961]: Showed exponential Diophantine relations are listable, reducing Hilbert’s 10th to finding a Diophantine representation of exponentiation. While their construction used higher degrees, our guard-gadget technique demonstrates degree $\leq 3$ suffices for universal representation over $\mathbb{N}$.
  2. 2.
    Jones [1982]: Achieved degree $\leq 4$ via direct MRDP reduction. Our guard-gadget technique (Definition 3.6) reduces this to degree $\leq 3$ by systematically shielding quadratic constraints.
  3. 3.
    Matiyasevich [1970]: Proved undecidability for degree $\leq 4$ (settling the “vanilla” Hilbert’s 10th Problem). Our result shows the degree barrier can be lowered by one, addressing a natural question raised by Jones [1982].
  4. 4.
    Hilbert’s Tenth Problem and its variants for rings, number fields, and related structures remain central questions in mathematics [Alpöge et al., 2025]. However, from a computational perspective, structures like $\mathbb{Z}$ reduce to $\mathbb{N}$ via standard encodings. Mathematical results, see [Poonen, 2009], concern definability complexity in strong foundational theories, which is orthogonal to our focus on bounded provability in weak arithmetics.
The remainder of the paper fills in the technical details:
  1. 1.
    Section 3 (Preliminaries): Establishes encoding tools ($\beta$-function, four-squares, Zeckendorf, guard-gadgets).
  2. 2.
    Section 4 (Construction): Builds the cubic system via proof sequence encoding, axiom testing, and modus ponens verification.
  3. 3.
    Section 5 (Aggregation): Merges the system via sum-of-squares and applies Jones factorization.
  4. 4.
    Section 6 (Computation): Provides algorithmic details for degree reduction.
  5. 5.
    Section 7 (Result): Proves soundness and completeness bidirectionally.
Each stage is constructed to be effective (algorithmically computable), polynomial-time (in the proof parameters), and degree-preserving (maintaining $\delta\leq 3$ after Stage 5).
@@ 3.1 The Gödel Beta-Function
Procedure 3.1 (Diophantine $\beta$-Function).
@@
In the completeness proof of Theorem 4.16, the witness construction algorithm (Exposition 4.9) must compute four-square decompositions for each remainder $r_{i}$ and its bound $(i+1)d-1-r_{i}$. By the above algorithm, this can be done in time $O(\log((i+1)d))$ per decomposition, yielding total time $O(N\cdot\log(\max_{i}f_{i}))$ for all $N$ proof lines. This is polynomial in the proof parameters, as required.
@@ 3.3 Combined Encoding
    Encode the sequence using $\beta(c,d,i)=f_{i}$ via constraint (4.47) and Lemma 3.2.
@@ 3.4 Summary of Polynomial Constraints
Constraint Expanded Equation $\boldsymbol{\delta}$ $\mathbb{N}$-Mechanism
Pre guard-wrapping
$\beta$-Division $c-q_{i}(1+(i+1)d)-r_{i}=0$ 2 Linear in $c,r_{i}$; quadratic in $q_{i},d$
Lower bound $r_{i}-a_{i}^{2}-b_{i}^{2}-c_{i}^{2}-e_{i}^{2}=0$ 2 Sum of squares $\geq 0$
Upper bound $(i+1)d-1-r_{i}-\alpha_{i}^{2}-\beta_{i}^{2}-\gamma_{i}^{2}-\delta_{i}^{2}=0$ 2 Sum of squares $\geq 0$
Boolean $d_{i,\kappa}-d_{i,\kappa}^{2}=0$ 2 Forces $d_{i,\kappa}\in\{0,1\}$
Non-adjacency $d_{i,\kappa}\cdot d_{i,\kappa+1}=0$ 2 Product vanishes
Zeckendorf sum $f_{i}-\sum_{\kappa=2}^{K}F_{\kappa}d_{i,\kappa}=0$ 1 Linear combination
Axiom membership $b_{ax,i}\cdot\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0$ 3 Boolean times sum-of-squares
MP arithmetic $f_{i}-f_{j}-f_{k}=0$ 1 Linear equality
Exactly-one $b_{ax,i}+\sum_{j,k<i}b_{mp,i,j,k}-1=0$ 1 Linear in booleans
Target check $f_{N}-g_{T}=0$ 1 Linear equality
Post guard-wrapping
$\beta$-Division guard $u_{\beta,i}\cdot(c-q_{i}(1+(i+1)d)-r_{i})=0$ 3 $u_{\beta,i}=1+v_{\beta,i}^{2}\geq 1$
Lower-bound guard $u_{L,i}\cdot(r_{i}-a_{i}^{2}-b_{i}^{2}-c_{i}^{2}-e_{i}^{2})=0$ 3 $u_{L,i}=1+v_{L,i}^{2}\geq 1$
Upper-bound guard $u_{U,i}\cdot((i+1)d-1-r_{i}-\alpha_{i}^{2}-\cdots)=0$ 3 $u_{U,i}=1+v_{U,i}^{2}\geq 1$
Boolean guard $u_{B,i,\kappa}\cdot(d_{i,\kappa}-d_{i,\kappa}^{2})=0$ 3 $u_{B,i,\kappa}=1+v_{B,i,\kappa}^{2}\geq 1$
Non-adjacency guard $u_{N,i,\kappa}\cdot(d_{i,\kappa}\cdot d_{i,\kappa+1})=0$ 3 $u_{N,i,\kappa}=1+v_{N,i,\kappa}^{2}\geq 1$
Zeckendorf guard $u_{Z,i}\cdot(f_{i}-\sum_{\kappa}F_{\kappa}d_{i,\kappa})=0$ 2 $u_{Z,i}=1+v_{Z,i}^{2}\geq 1$
Axiom guard (Already $\delta=3$ before wrapping) 3 Boolean activation
MP-sum guard $u_{mp,i,j,k}\cdot(f_{i}-f_{j}-f_{k})=0$ 2 $u_{mp,i,j,k}=1+v_{mp,i,j,k}^{2}\geq 1$
Exactly-one guard $u_{J,i}\cdot(b_{ax,i}+\sum_{j,k}b_{mp,i,j,k}-1)=0$ 2 $u_{J,i}=1+v_{J,i}^{2}\geq 1$
Target guard $\mathrm{T}\cdot(f_{N}-g_{T})=0$ 2 $\mathrm{T}=1+U^{2}\geq 1$
Positivity constraints
All guards $u_{\bullet}-1-v_{\bullet}^{2}=0$ 2 Forces $u_{\bullet}\geq 1$
Global multiplier $\mathrm{T}-1-U^{2}=0$ 2 Forces $\mathrm{T}\geq 1$
Table 2: Complete degree bookkeeping. Base constraints have $\delta\leq 3$; guard-wrapping preserves $\delta\leq 3$ for all main constraints, with guard positivity at $\delta=2$.